# Quantum Computing for Query Containmentof Conjunctive Queries

### v1

@@ 3.4. Using Constraints
A further optional step is to use constraints, i.e., to restrict the set $C$ of considered bit-matrices $\overline{x}$ to those bit-matrices that actually represent functions (i.e. where there is exactly one 1 per row). Such constraints can be imposed using an adapted QAOA version as explained in Section 2.2. This drastically reduces the search space, and additionally, it allows to simplify the polynomial $p(\overline{x})$ by completely dropping the term involving $p_{\textup{fct}}(\overline{x})$. Accordingly, for our choices $(p_{\textit{gen}},{d}_{\textit{gen}},C_{\textit{gen}})$ and $(p_{\textit{sim}},{d}_{\textit{sim}},C_{\textit{sim}})$, we get new constrained versions $(p^{c}_{\textit{gen}},{d}^{c}_{\textit{gen}},C^{c}_{\textit{gen}})$ and $(p^{c}_{\textit{sim}},{d}^{c}_{\textit{sim}},C^{c}_{\textit{sim}})$, which are defined as follows: $p^{c}_{\textit{gen}}(\overline{x}):=p_{\textup{(3)}}(\overline{x})$, $p^{c}_{\textit{sim}}(\overline{x}):=p^{\text{(s)}}_{\textup{(3)}}(\overline{x})$, ${d}^{c}_{\textit{gen}}:={d}_{\textit{gen}}=-|\textbf{T}_{2}|$, ${d}^{c}_{\textit{sim}}:={d}_{\textit{sim}}=-|\textbf{T}_{2}|$, and $C^{c}_{\textit{gen}}$ is the set of all bit-matrices $\overline{x}\in\{0,1\}^{B_{2}\times A_{1}}$ where for all $i\in B_{2}$ there is exactly one $j\in A_{1}$ with $x_{i,j}=1$, while $C^{c}_{\textit{sim}}$ is the set of all bit-matrices $\overline{x}\in\{0,1\}^{B^{\text{(s)}}_{2}\times A_{1}}$ where for all $i\in B^{\text{(s)}}_{2}$ there is exactly one $j\in A_{1}$ with $x_{i,j}=1$. Note that $|C^{c}_{\textit{gen}}|\ =\ (n_{1})^{n_{2}}\ \leqslant\ |C_{\textit{gen}}|=2^{n_{2}\cdot n_{1}}$ and $|C^{c}_{\textit{sim}}|\ =\ (n_{1})^{n^{\text{(s)}}_{2}}\ \leqslant\ |C_{\textit{sim}}|=2^{n^{\text{(s)}}_{2}\cdot n_{1}}$. Theorem 3.9 and 3.11 yield:
@@ 3.5. Identifying Trivial Cases
Let $(p,d,C)$ be any of the choices $(p_{\textit{gen}},{d}_{\textit{gen}},C_{\textit{gen}})$, $(p^{c}_{\textit{gen}},{d}^{c}_{\textit{gen}},C^{c}_{\textit{gen}})$, $(p_{\textit{sim}},{d}_{\textit{sim}},C_{\textit{sim}})$, $(p^{c}_{\textit{sim}},{d}^{c}_{\textit{sim}},C^{c}_{\textit{sim}})$. If the polynomial $p$ actually does not contain any binary variable (e.g. if all variables have been assigned during preparation and/or simplification), then $p(\overline{x})$ is the constant function $\ell$, for some integer $\ell$ (for short: $p(\overline{x})\equiv\ell$); see Appendix B for details. In case that $p(\overline{x})\equiv\ell$ for an integer $\ell$, we obtain from Theorem 3.9, Theorem 3.11 and Corollary 3.14 that $q_{1}\not\sqsubseteq q_{2}$ if $\ell\neq d$, and $q_{1}\sqsubseteq q_{2}$ if $\ell=d$; and in the latter case we can also construct a homomorphism from $q_{2}$ to $q_{1}$ that serves as a certificate which proves that $q_{1}\sqsubseteq q_{2}$.
@@
We continue Example 3.13 and consider the choice $(p_{\textit{sim}},{d}_{\textit{sim}},C_{\textit{sim}})$. Recall that $p_{\textit{sim}}(\overline{x})\equiv-2$ and ${d}_{\textit{sim}}=-2$. As explained above, this constitutes a trivial case where we can stop and output $q_{1}\sqsubseteq q_{2}$”. Inspecting the definition of $\overline{x}^{\textup{(s)}}$ yields a homomorphism from $q_{2}$ to $q_{1}$, namely the mapping $h\colon A_{2}\to A_{1}$ with $h(X_{2})=X_{1}$, $h(Y_{2})=Y_{1}$, $h(Z_{2})=Z_{1}$, and $h(W_{2})=\textsf{'actor'}$.
@@ 2. RQ2:
    Scalability: How does the probability of solving QC-CQ scale with the size of these problems?
@@ 5.1. Metrics
We determine how likely it is to successfully solve a QC-CQ problem and track the solution probability as the fraction of measurements (reads for annealing, shots for QAOA) for which an optimal solution is found. Here, we only consider positive query pairs, as the solution probability of negative pairs is always 100%. Small solution probabilities are not problematic; a single certifying homomorphism is evidence that containment holds. [^1] The probability $P^{*}$ of obtaining at least one solution grows exponentially with the number of measurements. For example, with a solution probability of only 10% and 20 measurements, the probability for a true positive reaches $P^{*}=1-0.9^{20}\approx 87.8\%$.
@@ 5.2. Experimental Setup
Query Pairs. We manually extracted query pairs from textbooks and benchmarks, which we refer to as T&B problems. Thus, these problems are community-authored. The textbook problems [^2] From (1) “Foundations of Databases” by Abiteboul, Hull, and Vianu (Abiteboul et al., 1995), (2) the online textbook “Database Theory” by Arenas, Barceló, Libkin, Martens, and Pieris (Arenas et al., 2022) (version from 2022-08-19), (3) the booklet “Answering Queries using Views” by Afrati, Chirkova, and Jagadish (Afrati et al., 2019), (4) lecture material by Koch taught at Saarland University in 2006 (shared in personal communication) and (5) lecture material by Schweikardt (Schweikardt, 2023). include the example from Sec. 2.1. We harvested various query equivalence or query containment benchmarks (uwd, 2025; idn, 2025; jen, 2025; Wudage Chekol et al., 2013; spa, 2025). Conjunctive SQL and SPARQL queries were manually converted to tableau form.
Figure 3(a) states the numbers of query pairs by their ground truth w.r.t. QC-CQ (positive/P or negative/N, derived manually). This includes problems that may later be identified as trivial. Figure 3(b) states the subsets of non-trivial problems that are evaluated in the quantum stage, via annealing (for a polynomial degree of up to 2) and QAOA (for any degree, with maximum occurring degree 4). The former are included in the latter. We state the original number of problems, as well as the number reduced by simplification.
@@
We designed two synthetic families of parameterized QC-CQ problems, based on the directed graph schema $\textbf{S}_{g}$ from Example 3.4. Figure 3(c) illustrates the archetypal building blocks. Family 2-cycle-to-$i$-chain contains pairs of a 2-cycle query and a query describing a chain of length $i$ for $i\in\mathbb{N}_{\geqslant 1}$. For illustration, revisit queries 2-cycle ($q_{\textit{2cy}}$) and 2-chain ($q_{\textit{2ch}}$) presented in Example 3.4. Family 2-chain-to-$i$-star, contains pairs of a 2-chain query and an $i$-star query for $i\in\mathbb{N}_{\geqslant 1}$. These problems can be solved with annealing and QAOA. They allow for small-stepped scaling in the number of binary variables, which here, map directly to qubits. Moreover, preparation or simplification do not reduce the number of binary variables. (While simplification has no effect, we will enable it by default in our experiments.) The search space is non-trivially exponential, yet we can manually derive two optimal (and only) solutions for each instance as ground truth (details in Appendix D).
Implementation. We implemented the workflow described in Section 4, using Python 3.10, qrisp 0.7.9, dwave-ocean-sdk 9.0.0, qiskit 2.2.1 and qiskit-aer 0.17.2. For experiments on the IQM device, we used Python 3.11, qrisp 0.7.17, and qiskit 2.1.2.
Backends. Table 1 provides details. We first describe the simulators. For annealing, we use the classical simulated annealer software SimulatedAnnealingSampler. The PathIntegralAnnealingSampler emulates the behavior of a quantum annealer. To simulate quantum circuit experiments, we use IBM Qiskit’s CPU-based Aer.
We use four quantum hardware devices. For quantum annealing, we use D-Wave’s Advantage System 6.4 and Advantage2 System 1.7. For QAOA, we use IQM Emerald and IBM Aachen.
@@
Standard configuration. Unless stated otherwise, we use a standard configuration. For easier comparison, we use the same configuration for experiments with simulators and quantum hardware. We deliberately keep the number of measurements low to ensure that successful outcomes cannot be attributed to random chance, and to meet the practical constraints of current hardware.
@@ 5.3. RQ1: Soundness
We first explore soundness of the workflow, and then focus on the quantum stage. We use simulators with standard configuration.
Soundness of end-to-end workflow. We process all positive T&B query pairs from Figure 3(a) with our 3-stage workflow. This includes problems that are trivial to start with, that become trivial after simplification, and that cannot be solved, as the degree of the polynomial is too high for annealing or where the number of qubits exceeds the simulator capabilities. We measure solution probabilities and therefore only consider the positive query pairs.
We distinguish two evaluation paths. First, we configure annealing as preferred solver and deviate from the workflow in Figure 2: If the polynomial degree is greater two, we return that the problem instance could not be solved (rather than defaulting to QAOA). Second, we configure QAOA as preferred solver.
@@
Figure 4. Soundness of 3-stage workflow, processing positive T&B query pairs (simulators). Showing fraction of problems where solution probability meets/exceeds a given threshold.
@@
We plot the fraction of query pairs whose solution probability meets/exceeds a given threshold. That is, for any value on the $x$-axis, the corresponding value on the $y$-axis indicates the fraction of pairs that achieved a solution probability greater than or equal to this threshold. To illustrate, we highlight three observations in Figure 4(a), Arrow (i) points out that a solution probability of 50% and greater is obtained with simulated annealing for about 80% of the query pairs (which accounts for almost all QUBO instances).
@@
The $y$-value at which the curve ends (indicated by Arrow (iii) at $x=1$) shows the share of problem instances that our approach solved with close to 100% solution probability. For reference, we plot the likelihood of solving the problem by chance (“random”), obtained by generating 500 bitvectors at random. [^3] For QAOA, we generate bit vectors that are valid: Each represents a matrix with one bit per row set to 1, to be comparable with constrained QAOA where this is enforced. This likelihood is considerably lower than the measured probabilities (apart from the corner cases of solution probabilities close to zero or one). The $y$-value at which the random curve ends (see Arrow (iv) at $x=1$) shows the fraction of problems that are trivial. Thus, the difference in fraction between Arrows (iii) and (iv) indicates the gains that can be contributed to our approach.
Throughout, we observe even higher solution probabilities with simplification enabled (even for guessing at random).
Discussion. With simplification, more input problems become trivial to start with, and many problems become easier to solve since the number of binary variables can often be reduced. Our results show that the simplification step is indeed very effective.
@@
Soundness of Annealing/QAOA step. We now focus on the quantum stage in isolation and only consider the nontrivial T&B problems (Figure 3(b)). We deliberately include the negative query pairs to check whether our objective function is implemented correctly.
@@ 5.4. RQ2: Scalability
However, 2-chain-to-$i$-star (Figure 5(b)) displays a strikingly different scaling behavior, and much higher solution probabilities. With simulated annealing (SA), smaller instances even yield lower solution probabilities than larger ones, and the curve quickly reaches 100%, so that each read returns an optimal solution. With QAOA, the solution probabilities remain higher compared to 2-cycle-to-$i$-chain, but there is a steep drop until all simulations exceed the timeout. To investigate the drop, we configured the “bigshot” configuration (purple dots in Figure 5(b)). Evidently, increasing the numbers of layers and shots improves the solution probabilities.
Discussion for RQ2. 2-cycle-to-$i$-chain scales as expected: the solution probability decreases with increasing problem size: Each additional variable increases the probability that the respective heuristic algorithm does not terminate in a global minimum. With QAOA, we exceed the capacity of the simulator sooner.
@@
Clearly, the scaling behavior is highly problem-specific. 2-cycle-to-$i$-chain displays the scaling behavior that can be expected based on state-of-the-art. Moreover, chains queries of length 20 an beyond constitute 20 self-joins, and may be considered not small. However, 2-cycle-to-$i$-star surprises with exceptionally good performance. This strongly motivates further research to investigate QC-CQ problems in terms of their energy landscapes.

### v2

@@ 3.4. Using Constraints
A further optional step is to use constraints, i.e., to restrict the set $C$ of considered bit-matrices $\overline{x}$ to those bit-matrices that actually represent functions (i.e. where there is exactly one 1 per row). Such constraints can be imposed using an adapted QAOA version as explained in Section 2.2. This drastically reduces the search space, and additionally, it allows to simplify the polynomial $p(\overline{x})$ by completely dropping the term involving $p_{\textup{fct}}(\overline{x})$. Accordingly, for our choices $(p_{\textit{gen}},{d}_{\textit{gen}},C_{\textit{gen}})$ and $(p_{\textit{sim}},{d}_{\textit{sim}},C_{\textit{sim}})$, we get new constrained versions $(p^{c}_{\textit{gen}},{d}^{c}_{\textit{gen}},C^{c}_{\textit{gen}})$ and $(p^{c}_{\textit{sim}},{d}^{c}_{\textit{sim}},C^{c}_{\textit{sim}})$, which are defined as follows: $p^{c}_{\textit{gen}}(\overline{x}):=p_{\textup{(3)}}(\overline{x})$, $p^{c}_{\textit{sim}}(\overline{x}):=p^{\text{(s)}}_{\textup{(3)}}(\overline{x})$, ${d}^{c}_{\textit{gen}}:={d}_{\textit{gen}}=-|\textbf{T}_{2}|$, ${d}^{c}_{\textit{sim}}:={d}_{\textit{sim}}=-|\textbf{T}_{2}|$, and $C^{c}_{\textit{gen}}$ is the set of all bit-matrices $\overline{x}\in\{0,1\}^{B_{2}\times A_{1}}$ where for all $i\in B_{2}$ there is exactly one $j\in A_{1}$ with $x_{i,j}=1$, while $C^{c}_{\textit{sim}}$ is defined analogously, by replacing $B_{2}$ with $B^{\text{(s)}}_{2}$. Note that $|C^{c}_{\textit{gen}}|\ =\ (n_{1})^{n_{2}}\ \leqslant\ |C_{\textit{gen}}|=2^{n_{2}\cdot n_{1}}$ and $|C^{c}_{\textit{sim}}|\ =\ (n_{1})^{n^{\text{(s)}}_{2}}\ \leqslant\ |C_{\textit{sim}}|=2^{n^{\text{(s)}}_{2}\cdot n_{1}}$. Theorems 3.9 and 3.11 yield:
@@ 3.5. Identifying Trivial Cases
Let $(p,d,C)$ be any of the choices $(p_{\textit{gen}},{d}_{\textit{gen}},C_{\textit{gen}})$, $(p^{c}_{\textit{gen}},{d}^{c}_{\textit{gen}},C^{c}_{\textit{gen}})$, $(p_{\textit{sim}},{d}_{\textit{sim}},C_{\textit{sim}})$, $(p^{c}_{\textit{sim}},{d}^{c}_{\textit{sim}},C^{c}_{\textit{sim}})$. If the polynomial $p$ actually does not contain any binary variable (e.g. if all variables have been assigned during preparation and/or simplification; see Appendix B for details), then $p(\overline{x})$ is the constant function $\ell$, for some integer $\ell$ (for short: $p(\overline{x})\equiv\ell$). In case $p(\overline{x})\equiv\ell$ for an integer $\ell$, we obtain from Theorems 3.9 and 3.11, and Corollary 3.14, that $q_{1}\not\sqsubseteq q_{2}$ if $\ell\neq d$, and $q_{1}\sqsubseteq q_{2}$ if $\ell=d$; in the latter case we can construct a homomorphism from $q_{2}$ to $q_{1}$ as a certificate proving $q_{1}\sqsubseteq q_{2}$.
@@
We continue Example 3.13 and consider the choice $(p_{\textit{sim}},{d}_{\textit{sim}},C_{\textit{sim}})$. Recall that $p_{\textit{sim}}(\overline{x})\equiv-2$ and ${d}_{\textit{sim}}=-2$. As explained above, this constitutes a trivial case where we can stop and output $q_{1}\sqsubseteq q_{2}$”. Inspecting the definition of $\overline{x}^{\textup{(s)}}$ yields a homomorphism from $q_{2}$ to $q_{1}$, namely the mapping $h\colon A_{2}\to A_{1}$ with $h(X_{2})=X_{1}$, $h(Y_{2})=Y_{1}$, $h(Z_{2})=Z_{1}$, and $h(W_{2})=\textsf{\textquotesingle actor\textquotesingle}$.
@@ 4. System and Workflow
Performing several measurements, we obtain a set of bitvectors, where each encodes a bit-matrix.
@@ 2. RQ2:
    Scalability: How does the probability of solving QC-CQ problems scale with the size of these problems?
@@ 5.1. Metrics
We determine how likely it is to successfully solve a QC-CQ problem and track the solution probability as the fraction of measurements (reads for annealing, shots for QAOA) for which an optimal solution is found. Here, we only consider positive query pairs, as the solution probability of negative pairs is always 100%. As long as the solution probability is non-zero, we find a certifying homomorphism which proves that containment holds. [^1] The probability $P^{*}$ of obtaining at least one solution grows exponentially with the number of measurements. For example, with a solution probability of only 10% and 20 measurements, the probability for a true positive reaches $P^{*}=1-0.9^{20}\approx 87.8\%$.
@@ 5.2. Experimental Setup
Query Pairs. We manually extracted query pairs from textbooks and benchmarks. Thus, these “T&B problems” are community-authored. The textbook problems [^2] From (1) “Foundations of Databases” by Abiteboul, Hull, and Vianu (Abiteboul et al., 1995), (2) the online textbook “Database Theory” by Arenas, Barceló, Libkin, Martens, and Pieris (Arenas et al., 2022) (version from 2022-08-19), (3) the booklet “Answering Queries using Views” by Afrati, Chirkova, and Jagadish (Afrati et al., 2019), (4) lecture material by Koch taught at Saarland University in 2006 (shared in personal communication), and (5) lecture material by Schweikardt (Schweikardt, 2023). include the example from Sec. 2.1. We harvested various query equivalence or query containment benchmarks (uwd, 2025; idn, 2025; jen, 2025; Wudage Chekol et al., 2013; spa, 2025). Conjunctive SQL and SPARQL queries were manually converted to tableau form.
Figure 3(a) states the numbers of query pairs by their ground truth w.r.t. QC-CQ (positive/P or negative/N, derived manually). This includes problems that may later be identified as trivial. Figure 3(b) states the subsets of non-trivial problems that are evaluated in the quantum stage, via annealing (for a polynomial of degree up to 2) and QAOA (for any degree, with maximum occurring degree 4). The former are included in the latter. We state the original number of problems, as well as the reduced number, after simplification.
@@
We designed two synthetic families of parameterized QC-CQ problems, based on the directed graph schema $\textbf{S}_{g}$ from Example 3.4. Figure 3(c) illustrates the archetypal building blocks. Family 2-cycle-to-$i$-chain contains pairs of a 2-cycle query and a query describing a chain of length $i$ for $i\in\mathbb{N}_{\geqslant 1}$. For illustration, revisit the queries 2-cycle ($q_{\textit{2cy}}$) and 2-chain ($q_{\textit{2ch}}$) presented in Example 3.4. Family 2-chain-to-$i$-star, contains pairs of a 2-chain query and an $i$-star query for $i\in\mathbb{N}_{\geqslant 1}$. All these problems can be solved with annealing and QAOA. They allow for small-stepped scaling in the number of binary variables, which here, map directly to qubits. Moreover, preparation or simplification do not reduce the number of binary variables. (While simplification has no effect, we will enable it by default in our experiments.) The search space is non-trivially exponential, yet we can manually derive two optimal (and only) solutions for each instance as ground truth (details in Appendix D).
Implementation. We implemented the workflow described in Section 4, utilizing Eclipse Qrisp for gate-based QAOA across multiple backends and D-Wave libraries for annealing. Specifically, we use Python 3.10, qrisp 0.7.9, dwave-ocean-sdk 9.0.0, qiskit 2.2.1 and qiskit-aer 0.17.2. To access IQM quantum hardware, we used Python 3.11, qrisp 0.7.17, and qiskit 2.1.2.
Backends. We first describe the simulators. For annealing, we use the classical simulated annealer software SimulatedAnnealingSampler. The PathIntegralAnnealingSampler emulates the behavior of a quantum annealer. To simulate quantum circuit experiments, we use IBM Qiskit’s CPU-based Aer. We use four quantum hardware devices. For quantum annealing, we use D-Wave’s Advantage System 6.4 and Advantage2 System 1.7. For QAOA, we use IQM Emerald and IBM Aachen. Table 1 provides details on all solvers.
@@
Standard configuration. Unless stated otherwise, we use a standard configuration. For easier comparison, we use the same configuration for experiments with simulators and quantum hardware. We deliberately keep the number of measurements low to ensure that successful outcomes cannot be attributed to random chance and to meet the practical limitations of current hardware.
@@ 5.3. RQ1: Soundness
We first explore soundness of the entire workflow and then of the quantum stage. We use simulators with standard configuration.
Soundness of end-to-end workflow. We process all positive T&B query pairs from Figure 3(a) with our 3-stage workflow. This includes problems that are trivial to start with, that become trivial after simplification, and that cannot be solved, as the degree of the polynomial is too high for annealing or where the number of qubits exceeds the simulator capabilities. We measure solution probabilities, and therefore only consider the positive query pairs.
We distinguish two evaluation paths. For the first, we configure annealing as preferred solver and deviate from the workflow in Figure 2: For polynomial degree greater two, we return that the problem instance could not be solved (rather than defaulting to QAOA). For the second, we configure QAOA as preferred solver.
@@
Figure 4. Soundness of 3-stage workflow of positive T&B query pairs (on simulators). Showing fraction of problems where solution probability meets/exceeds a given threshold.
@@
We plot the fraction of query pairs whose solution probability meets/exceeds a given threshold. That is, for any value on the $x$-axis, the corresponding value on the $y$-axis indicates the fraction of pairs that achieved a solution probability greater than or equal to this threshold. To illustrate, we highlight three observations in Figure 4(a). Arrow (i) points out that a solution probability of 50% and greater is obtained with simulated annealing for about 80% of the query pairs (which accounts for almost all QUBO instances).
@@
The $y$-value at which the curve ends (see Arrow (iii) at $x=1$) shows the share of problem instances solved with close to 100% solution probability. For reference, we plot the likelihood of solving the problem by chance (“random”), obtained by generating 500 bitvectors at random. For QAOA, we generate bitvectors that are valid: Each represents a matrix with one bit per row set to 1, to be comparable with constrained QAOA where this is enforced.
This likelihood is considerably lower than the measured probabilities (apart from the corner cases of solution probabilities close to zero or one). The $y$-value at which the random curve ends (Arrow (iv) at $x=1$) shows the fraction of problems that are trivial. Thus, the difference in fraction between Arrows (iii) and (iv) is the gain that can be contributed to our approach.
Throughout, we observe even higher solution probabilities when simplification is enabled (even for guessing at random).
Discussion. With simplification, more input problems become trivial to start with, and many problems become easier to solve, since the number of binary variables can often be reduced. Our results show that the simplification step is indeed very effective.
@@
Soundness of quantum stage. We now focus on the quantum stage in isolation and only consider the nontrivial T&B problems (Figure 3(b)). We deliberately include the negative query pairs to check whether our objective function is implemented correctly.
@@ 5.4. RQ2: Scalability
However, 2-chain-to-$i$-star (Figure 5(b)) displays a strikingly different scaling behavior, and much higher solution probabilities. With simulated annealing (SA), smaller instances even yield lower solution probabilities than larger ones, and the curve quickly reaches 100%, i.e., each read returns an optimal solution. With QAOA (AerSim), the solution probabilities remain higher compared to 2-cycle-to-$i$-chain, but there is a steep drop until all simulations exceed the timeout. To investigate, we compare with the “bigshot” configuration (purple dots in Figure 5(b)). Evidently, increasing the number of layers, shots, and iterations improves the solution probabilities.
Discussion. 2-cycle-to-$i$-chain scales as expected: the solution probability decreases with increasing problem size: Each additional variable increases the probability that the respective heuristic algorithm does not terminate in a global minimum. With QAOA, we exceed the capacity of the simulator sooner.
@@
Clearly, the scaling behavior is problem-specific. 2-cycle-to-$i$-chain scales as can be expected based on state-of-the-art, and chain queries of length 20 are no longer small queries. However, 2-cycle-to-$i$-star surprises with exceptionally good performance. This strongly motivates further research to investigate QC-CQ problems in terms of their energy landscapes.